A bullet of mass 10grms is travelling horizontally with a velocity of 150 m/sec strikes a stationary wooden block and comes to arest after 0.03 seconds. Calculate the distance of prnitration of the bullet into the wooden block.


please help!


 

Grade:9

2 Answers

Radhika Batra
13 years ago

Dear neel,

Given that :

Initial velocity (u) = 150 m/s.

Final velocity (v)  =  0 m/s.

Time taken for the body to come at rest (t)= 0.03 sec.

Now according to newton's first law:

 v = u - at   -------------------------------------------  eq 1.

so from this equation value of :  a = (u/t).

Now from the euation we have :

s = ut - 1/2at2 ----------------------------------------- eq 2.

put the values of u, t, a in the 2nd equation.

after solving we will get that :

s = 4.5m.

Thus the bullet will  penetrate 4.5 meters inside the wodden block.

 

askIITians.com provides online iit jee courses and IIT JEE Test Series with IITians. Click here to get free online test series and check your status timely or you can join us as our registered user for getting best iit jee study material or iit jee test series.

Click here to find IIT JEE Packages or course details with askIITians.

for further details visit the folowing links:

http://www.askiitians.com//iit-jee-online-coaching.aspx

http://www.askiitians.com/iit-jee-test-series.aspx

http://www.askiitians.com/iit-study-material/

Askiitians provide you all the best possible information, feel free to ask.

all the best.

thanks and regards.

sidharth G
13 years ago

mass of bullet = 0.01 kg

initial vel ,u = 150 m/s

final vel , v= 0

time= 0.03s

distance = s

v=u+at

v2=u2+2as

use these eq to find a(accelaration) and s respectively .u will get the answer